A body of mass 100 gram, tied at the end of a string of length 3 m rotates in a vertical circle and is just able to complete the circle. If the tension in the string at its lowest point is 3.7 N, then its angular velocity will be ______ (g = 10 m/s2)

Grade:12

1 Answers

Sadaf
5 years ago
As the tension is at lowest piont so tension will be upwards weight downward and centripetal force upward
So
T-mg=mv^2/r
3.7-0.1*10= 0.1v^2/r
V=9m/s
We know
V=wr
9=w*3
W=3
That is angular velocity
 
V= wr
9=